OpenStudy (czarluc):

(x^2-7x+12)^x^2-3x+2=1 solve for x

OpenStudy (anonymous):

Is the question (x^2-7x+12)^(x^2-3x+2)=1 ? .@czarluc

OpenStudy (czarluc):

the x^2-3x+2 is the exponent

OpenStudy (czarluc):

yes you are right princer

OpenStudy (anonymous):

This given expression holds only if x^2-3x+2=0 that is if (x-1)(x-2)=0 that is if x=1,2 or if x^2-7x+12=1 i.e is if x^2-7x+11=0 i.e. if x={7+square root(5)}/2 or x={7-square root(5)}/2 @czarluc

OpenStudy (czarluc):

oh ok I got it! thanks!